### v2

# inkscapelatex=false

Abstract
Sequential activation of place-tuned neurons in an animal during navigation is typically interpreted as reflecting the sequence of input from adjacent positions along the trajectory. More recent theories about such place cells suggest sequences arise from abstract cognitive objectives like planning. Here, we propose a mechanistic and parsimonious interpretation to complement these ideas: hippocampal sequences arise from intrinsic recurrent circuitry that propagates activity without readily available input, acting as a temporal memory buffer for extremely sparse inputs. We implement a minimal sequence generator inspired by neurobiology and pair it with an actor–critic learner for egocentric visual navigation. Our agent reliably solves a continuous maze without explicit geometric cues, with performance depending on the length of the recurrent sequence. Crucially, the model outperforms LSTM cores under sparse input conditions (16 channels, $\sim 2.5\%$ activity), but not under dense input, revealing a strong interaction between representational sparsity and memory architecture. In contrast to LSTM agents, hidden sequence units develop localized place fields, distance-dependent spatial kernels, and task-dependent remapping, while inputs orthogonalize and spatial information increases across layers. These phenomena align with neurobiological data and are causal to performance. Together, our results show that sparse input synergizes with sequence-generating dynamics, providing both a mechanistic account of place cell sequences in the mammalian hippocampus and a simple inductive bias for reinforcement learning based on sparse egocentric inputs in navigation tasks.
Keywords Hippocampus $\cdot$ Visual navigation $\cdot$ Place cell $\cdot$ Theta rhythm $\cdot$ Reinforcement learning
1 Introduction
The hippocampus is a brain region that is essential for episodic memory and spatial navigation. Damage impairs memory formation in humans (Scoville and Milner, 2000), and rodent recordings revealed place cells tuned to specific locations (O’Keefe and Dostrovsky, 1971). These findings support the idea that spatial anchors structure episodic memory (Aronowitz and Nadel, 2023), reinforced by observations of replay sequences correlated to planning (Foster and Wilson, 2007; Kay et al., 2020).
Recent computational models have recast these findings through modern machine learning. Successor representations interpret place cells as predictive states (Stachenfeld et al., 2017; Mattar and Daw, 2018); reservoir models emphasize pre-existing dynamics (Leibold, 2020); probabilistic approaches infer place fields as latent states from sequential sensory experiences(Raju et al., 2024); and self-supervised methods learn spatial tuning from movement statistics (Wang et al., 2024). While these approaches reproduce place-like activity and even sequential patterns, they rarely address explicitly where hippocampal sequences originate.
We propose a parsimonious account: hippocampal theta sequences, that are experimentally observed during navigation Foster and Wilson (2007), arise from intrinsic recurrent circuitry Fenton (2024) in CA3 that can propagate activity over long timescales in the absence of input (Fig. 1.) The CA3 sequence generator receives sparse inputs from dentate gyrus (DG) to yield localized spatial codes in CA3 that support navigation (Leibold, 2020, 2022). The sparse-input regime is not incidental but central: it reflects ecological contraints of navigation often guided by only a few reliable landmarks amid abundant noisy sensory input; the biology of DG granule cells, which fire at extremely low mean rates; and the computational advantages of high-capacity, low-interference codes that promote high memory capacity, compositionality and generalization.
From a machine learning perspective, our architecture resonates ideas from dynamical systems models for long-range memory such as LSTMs Hochreiter and Schmidhuber (1997) and echo state machines Jaeger and Haas (2004). More recently, state-space models (SSMs) and linear RNNs have been suggested that preserve information by expanding inputs into high-dimensional representations before compressing them with shallow nonlinear readouts (Fu et al., 2022; Gu et al., 2020, 2021). Our model extends these principles with a DG sparsification stage that generates a low-activity code optimized for the CA3 recurrent network, whose intrinsic sequence dynamics are particularly effective under sparse input conditions, offering a rich set of features for downstream policy learning from compact visual features.
To test the hypothesis that sparse inputs to a long-horizon sequence-generating dynamical system are sufficient to give rise to a representation allowing successful navigation, we implement an agent with a DG-like sparsification module, a recurrent sequence generator (CA3 proxy), and an actor–critic learner for egocentric navigation. We show that sequence generation and sparse input synergize, outperforming LSTMs of comparable size in the sparse-input regime, while LSTMs remain competitive under dense input. Moreover, place-field like neural tuning, DG orthogonalization, and task-dependent remapping emerge naturally during training.
These results suggest that a sequence-based reservoir, inspired by CA3, is well suited for constructing spatial representations from sparse low-bandwidth inputs. The synergy between sparse coding and intrinsic sequence dynamics thus offers both a mechanistic explanation for hippocampal sequences and a simple inductive bias for reinforcement learning in navigation tasks.
Figure 1: Theta sequences. A Illustration of theta sequences observed in rodent hippocampus. In each theta cycle, R=3 neurons are activated and the activation propagates in a sequence of $\ell$=5 neurons over L=3 theta cycles (cf. eq. 2). B Current baseline understanding of the theta sequences, driven by sequential inputs despite recurrent connections in CA3. Left: spatial tunings of the sequentially activated cells. C Intrinsic theta sequence hypothesis, a parsimonious account where the recurrent connections support generating long horizon sequential activity without sequential external inputs.
2 Methods
Figure 2: Model summary. A Virtual environment ($19\times 19$ tiles) were constructed with walls randomly placed on 15 % of the tiles. Wall layouts for kept fixed for repeated trials, with an invisible goal (red star) near to the bottom right unless otherwise mentioned. At each episode, the agent was initially placed at a random location at least 5 tiles away from the goal. B The agent receives a first person perspective vision input that is processed via a visual encoder (ResNet, matching the SOTA in deepmind lab environment (Espeholt et al., 2018), pretrained and fixed in our experiments; cf. A3). C These output was linearly transformed to F features, and then sparsified using batch normalization and high thresholding ($\tau=2.43$), such that the percentage of activation ($\sim 2.5\%$) matches the sparse activity of neurons in the hippocampal DG that project to CA3 area. The modules in white are hard coded or fixed during training. The modules in blue (fully connected layers, FC) are trained.
2.1 Virtual Environment
We consider an agent navigating a continuous environment with sparse obstacles and uniform visual textures. The 3-d virtual environment was simulated using Deepmind lab (Beattie et al., 2016) with 60 fps, each action is repeated for 8 frames until the next observation. Visual observations are captured at half the wall height, with FOV of $90\degree$ and resolution of $92\times 76$ pixels in RGB channels. The map has $19\times 19$ squares (excluding the confining walls) with obstacle walls randomly placed on 15 % of the tiles. The wall layouts were kept fixed for repeated episodes unless otherwise stated. The environment was designed such that spatial relations between locations cannot be trivially inferred from the similarity of their corresponding visual features and that it allows multiple possible trajectories from a given location to the target.
In each episode, the agent was initially placed in a random location at least 5 units from the goal. An episode is finished when the agent reaches the goal location or a maximum of 7200 frames (with action repeat=8 corresponding to a maximum of 900 action steps). For the map displayed in Figure 2, the reward was placed near the bottom right corner and then moved to the lower left corner to test generalization. Since our navigation task does not require complex actions, we reduced the action space used in IMPALA (Espeholt et al., 2018), to facilitate training speed (Table A2).
2.2 Visual processing
The visual input is intended to mimic visual cortical preprocessing, extracting general-purpose visual features for the hippocampus. We pretrained a ResNet encoder (He et al., 2015) in combination with an LSTM core as in Espeholt et al. (2018) and kept it fixed when training our hippocampus model. We also tested a variant using the layer 2 output from ResNet pretrained on Imagenet. The model produces similar performance (Fig. A1).
2.3 Hippocampus Model
Dentate gyrus as a sparsification module.
The main cortical input to the hippocampus is routed through the dentate gyrus (DG) (Amaral et al., 2007). DG activity is characteristically sparse, with only a small fraction of granule cells active in a given environment (Henze et al., 2000; Leutgeb and Leutgeb, 2007). We model the DG as a linear projection of visual features followed by batch normalization (running estimates of mean and variance, retention rate $0.95$ per minibatch) and a high activation threshold to match the sparsity of activity ($\sim 2.5\%$).
CA3 as a sequence-generating shift register.
We model CA3 as a linear RNN shift register that propagates inputs as theta sequences (Fig. 1A) and, for this paper, we keep it fixed to isolate the effects of intrinsic sequence generation from effects potentially induced by recurrent plasticity. The DG provides input $u_{t}\in\mathbb{R}^{F}$ over $F$ features. Each feature is assigned a dedicated prewired sequence of length $\ell$, so the CA3 state is $X_{t}\in\mathbb{R}^{F\ell}$. Motivated by hippocampal theta sequences (Dragoi and Buzsáki, 2006; Foster and Wilson, 2007; Leibold, 2020), we parameterize the total number of sequence units as $\ell=L+(R-1)$, where $L$ is the number of theta cycles and $R$ the number of active units per cycle.
Single-feature dynamics ($F=1$).
Let $x_{t}\in\mathbb{R}^{\ell}$ denote the CA3 state for a single feature and $u_{t}\in\mathbb{R}$ the corresponding DG input. Sequence propagation is
$x_{t+1}\;=\;S\,x_{t}\;+\;J\,u_{t},$ (1)
where $S\in\mathbb{R}^{\ell\times\ell}$ is the shift operator and $J\in\mathbb{R}^{\ell\times 1}$ injects the input into the first $R$ slots:
$S=\begin{bmatrix}0&0&\dots&0\\ 1&0&\dots&0\\ &\ddots&\ddots&\vdots\\ 0&\dots&1&0\end{bmatrix}_{\ell\times\ell},\qquad J=\bigl[\underbrace{1,1,\dots,1}_{R\text{ times}}\;,\underbrace{0,0,\dots,0}_{L-1\text{ times}}\bigr]^{T}\in\mathbb{R}^{\ell\times 1}.$ (2)
Thus a transient input $u_{t}$ creates activity in the first $R$ positions which is then shifted one step per timestep along the length-$\ell$ register.
Multiple features ($F>1$).
Each feature dimension evolves independently under the same dynamics. Stacking all $F$ sequences with Kronecker expansion gives the block-structured update
$A\;=\;I_{F}\otimes S,\qquad B\;=\;I_{F}\otimes J,$ (3)
and the full CA3 dynamics
$X_{t+1}\;=\;A\,X_{t}\;+\;B\,u_{t},$ (4)
with fixed recurrent matrix $A$ and input matrix $B$.
2.4 Actor Critic Network
The activity of hippocampal units is fed into a decoder module with two linear/fully connected (FC) layers with ReLU activation. The actions and value are computed as the linear readout from the decoder module. We trained the agent with a standard advantage actor–critic objective (policy-gradient + value-baseline + entropy regularization) as implemented in Sample Factory (Petrenko et al., 2020; Espeholt et al., 2018).
Figure 3: Training performance A Agents with different sequence length L and repetition number R. Line and shaded area are mean and s.e.m. across 6 random seeds. Len: number of frames to reach the goal location. Stable success: the rate of having 100 consecutive successful episodes. Metrics were Gaussian smoothed, $\sigma=6\times 10^{6}$ frames. B The best performing agent with L=64 and R=8 across all seeds was tested for its transfer learning. New Rew.: new reward location at the lower left corner. new map: a randomly generated map with the same statistics. block path: new walls are added to block paths to the reward, while the previous blocking walls are removed. C and D Performance of agents with different recurrent modules and inputs. C: sparse input. D: dense input where batch-normalization and high thresholding was removed. CA3: our CA3 model with L=64 and R=8. RandRNN: randomly initialized fixed RNN of the same state size. SSM_LegS: fixed SSM HiPPO-LegS from Gu et al. (2020) with the same state size. LSTM: trainable LSTM with matching number of parameters.
3 Results
3.1 Behavioral Performance
Training the naive agent (L=64, R=8) on one fixed reward location in the maze exhibits robust performance measures after about 350 million frames (Fig. 3A). Thus our model effectively maintains information about sparse inputs in its trajectories of the RNN-like dynamics across time like in a reservoir (Jaeger and Haas, 2004). This view is further supported by our simulations with reduced sequence length showing inferior performance (Fig. 3 A). In the most extreme case ($L=1,R=1$) without sequences, essentially a pure feedforward architecture corresponding to a brain with DG output bypassing CA3, the agent did not achieve robust behavior. While agents with reduced sequence lengths ($L=1,16,R=8$) still can express reasonable success rates and trajectory lengths, their behavior is considerably more unstable as evidenced by the low fraction of consecutive successes (Fig. 3A). Transfer learning for new reward location and blocked paths, however, requires only about 50 million frames, and transfer learning to a new map requires about 150 million frames. These indicate that the agent was able to form a generalizing representation of the map and task (Fig. 3B).
The agent equipped with a sequence-generating module (CA3; eq.4) learns faster under sparse DG input than an agent in which CA3 is replaced by an LSTM core, resembling the SOTA architecture of Hessel et al. (2019) on the DMLab-30 benchmark (Beattie et al., 2016), from which our environment is adapted. Our CA3 module also outperforms state-space model HiPPO-LegS (Gu et al., 2020) and, as an additional control, randomly initialized RNNs, indicating that its theta-sequence dynamics provide a distinct advantage (Fig.3C).
Crucially, this advantage is confined to the sparse-input regime. Under dense input, LSTMs perform better—consistent with prior reports (Hessel et al., 2019)—and our CA3 module performs worse than HiPPO-LegS and random RNNs (Fig. 3D). We also tested other variants from Gu et al. (2020) and they performed similar to HiPPO-LegS (Tab. 1, Fig. A3. These results highlight a specific synergy between sparse representations and intrinsic sequence generator.
Figure 4: Evolution of occupancy over the course of learning. Color represents a normalized measure of time the agent spent at a location. Mean head directions at a location were visualized in arrows. From epoch 5 onwards, the agent has a preference to reach the bottom right corner (independent of the random starting location) and proceed to the goal location (red star) from there.
3.2 Occupancy
We divided the training into 6 epochs and evaluated the behavior and learned representations at these checkpoints. The agent develops a stable trajectory after 4 to 5 epochs. Between Epoch 4 and Epoch 5, the agent learned to get around the obstacle in the upper part of the right edge and started to spend more time at the lower right corner before reaching the reward site (Fig. 4). The agent appears to develop a strategy of visiting locations with salient input/landmarks and converging from different starting locations to the habituated paths, even after switching to a new reward location. This goes in line with the typical strategies employed by animals in familiar enviroments Gibson et al. (2013). In comparison, the LSTM agent under dense sensory input has fewer converging trajectories before reaching the reward (Fig. A8, likely due to the goal information being more readily available in the visual input, thereby implementing a strategy more related to visual search.
Table 1: Performance comparison across architectures. Training steps indicate the number of environment frames (in millions) required to reach 80% success . “✗” indicates that the threshold was not reached in all random seeds within 350M frames.
Input CA3 Random RNN HiPPO-LegS HiPPO-LegT HiPPO-LagT LSTM
[…]
Figure 5: Left: spatial tuning of DG and CA3 units from epoch 6. Pixel coordinates correspond to environment (black crosses correspond to walls). Each row shows the CA3 units ordered by their positions in the activity sequence. We selected 4 out of the 16 feature sequences for visualization. Spatial Information (SI): bits per time step. Right: spatial tuning of randomly selected Decoder layer 1 units from epoch 6.
Figure 6: Correlation (color code) of the DG activity maps evoked by the 16 (rows and columns) visual features. Throughout training, the DG input to CA3 became increasingly orthogonal.
3.3 Place Field Analysis
In order to understand how the agent executes a successful strategy and whether the representations in the hippocampus-inspired model would also align with the known physiology of place cells, we further computed activity maps of the units for the different layers of the model network. For the input units (DG), we (by construction) observe sparse activation that is distributed across few locations in the map (Fig. 5). More interestingly, learning the weights from the visual encoder to DG leads to orthogonalization, such that DG population activity over time develops a unique representation of individual locations (Fig. 6). Units positioned later in CA3 sequences (e.g. columns no.48 and no. 64 in Fig. 5) show broader spatial tuning, in line with experimental reports on the CA3 spatial code being less sparse than in DG (Jung and McNaughton, 1993). In contrast, the input to LSTM agents did not show orthogonalization of the learned input projection (Fig. A11) and the spatial tuning of the LSTM units are disperse across the entire map, distinct from localized place-cell-like tunings (Fig. A9.
Figure 7: SI of units in CA3 as a function of their order in a sequence and its evolution across epochs. Each sequence is indicated by a unique color across epochs.
3.4 Spatial Information Analysis
Qualitative inspection of the spatial tunings in Fig. 5 is also quantitatively supported by the distributions of spatial information (SI) of the neural activity maps: CA3 units activated $\sim 16$ steps after the input have larger SI (Fig. 7). This corroborates our training results that agents with sequence length $L=16$ are able to acquire the task reasonably well (Fig. 3). Units activated later in a sequence exhibit smaller spatial information (Fig. 7).
Tracking the distribution of SI rates across training epochs, we observe a growing long tail indicative for the development of a neural space representation throughout all layers (Fig. A6). Furthermore, SI also increases with increasing hierarchical level of the network layer until the first layer of the decoder network.
To test whether SI is causal to behavior, we selectively permute the output weights of Decoder layer 1 units. When the permutation was done on the 32 units with the lowest SI, the performance was not impaired with respect to success rate and trajectory length (100% and 1065 frames). When the permutation was done on the 32 units with the highest SI, the success rate dropped by 4.9% and the average trajectory length increased from 1065 frame to 2794 frames.
3.5 Population-level Representations
We next examined population-level representations using mean population vector (PV) correlations, defined as the Pearson correlation between two activity vectors separated by a spatial displacement $\Delta x,\Delta y$. This measure can be interpreted as a spatial kernel learned by the network (Fig. 8). Kernels from all layers (DG, CA3, and Decoder layers) exhibit some dependence on distance throughout learning. However, CA3 kernels are smoother than the DG kernels. Both DG and CA3 show lower correlation values compared to the actor–critic layers. Notably, the first Decoder layer (Decoder1, biologically approximately equivalent to area CA1) develops the most pronounced spatial tuning (Fig. 8), consistent with the single-unit selectivity observed in Fig. A6. After the reward location was changed (“new”), kernels became less sharply defined, indicating that the representations are disrupted although behavior is adapted. Most layers in the agent with LSTM core and dense input did not show spatial kernels with strong displacement dependency. Only the LSTM output units showed a gradually refined spatial kernel during learning but it is strongly non-isotropic (Fig. A12).
Figure 8: Kernel representation of distance. A Mean Pearson correlation (color code) between two population activity vectors as a function of their spatial displacement. Rows correspond to different layers (brain regions) of the network, columns to learning stages (epochs). B Left: summary of population vector correlation as a function of distance. Right: summary of population vector correlation as a function of orientation. Radial coordinate range: (-0.5,1.1).
Over training, a stable place code gradually emerges (Fig. A7), with DG similarities amplified in downstream layers. Interestingly, the similarity between fully trained networks (epoch 6) and the novel reward condition (“new”) is higher than between the naive (epoch 0) and trained (epoch 6) networks, suggesting that the agent acquires generalizable knowledge about the arena’s spatial layout.
4 Discussion
In this study, we presented a minimal model of the hippocampus that enables navigation in a vision-based virtual environment and reproduces key phenomena of spatial representations observed in the mammalian hippocampus.
The cornerstone of our model—the sequential connectivity in CA3—is biologically plausible and inspired by established findings on theta sequence firing Dragoi and Buzsáki (2006); Foster and Wilson (2007). By constraining CA3 recurrent connectivity as a dynamical reservoir and limiting training to input and output weights via reinforcement learning, we clearly isolate the effect of this theta sequence generator.
Agents equipped with the CA3 sequence module achieved superior navigation from sparse input (analogous to DG activity) compared to LSTMs of similar complexity or SSM-based cores. Crucially, this advantage was regime-dependent: under sparse input, CA3 dynamics and DG sparsification synergized to support robust navigation, whereas under dense input, conventional recurrent cores such as LSTMs performed better. Performance also degraded substantially when CA3 sequence length was ablated or shortened (Fig. 3), highlighting the functional role of theta sequence generator.
Analysis of neural activities revealed clear parallels with experimental observations of place cell properties, including robust place field formation (Fig. 5), population-level encoding of spatial distance (Fig.8), progressive orthogonalization of DG outputs (Fig. 6), and dynamic remapping triggered by changes in reward locations (Fig.A7), consistent with experimental findings Leutgeb and Leutgeb (2007); Fenton (2024).
Interpretation.
The CA3 module expands sparse DG codes into a temporally smoothed canonical basis set, providing long-horizon history without the indiscriminate feature mixing of fully connected RNNs. This is especially beneficial under sparse input, where immediate sensory signals are limited and long-range context is critical for policy learning.
At first sight, sparsifying input to a recurrent core seems purely detrimental—leaving the agent “blind” most of the time. Yet it also filters out noisy, non-informative cues, making each supra-threshold input more reliably tuned to restricted regions of space. Through sequence propagation and policy learning, representations of other locations can then be inferred and stabilized. This mechanism is consistent with habitual trajectories: neuronal activity late in a sequence only remains informative if the policy converges onto consistent paths.
In contrast, with dense input streams, conventional recurrent cores such as LSTMs performed better, consistent with their ability to rapidly integrate continuous signals. Our CA3 model, however, performed worse under dense input. A likely reason is that its long-range memory buffer provides little benefit when input is already abundant, and temporal smoothing obscures informative short-term variations.
Overall, these results suggest that different recurrent architectures are suited to different sensory regimes: sparse coding naturally complements sequential expansion, whereas dense input favors mixing-oriented recurrent dynamics. The agents’ behaviors also diverged, resembling memory-driven navigation in our model versus more sensory-driven “visual search” navigation with LSTMs.
Biological and computational relevance.
We proposed a parsimonious account of hippocampal theta sequences: they can be intrinsically maintained without requiring external input. While actual theta sequences in biological systems likely arise from multiple factors and vary across contexts Chance (2012); Schlesiger et al. (2015); Ahmadi et al. (2025), our abstraction nonetheless reproduced hallmark hippocampal phenomena, including robust place fields (Fig. 5), progressive DG orthogonalization (Fig. 6), distance-dependent population kernels (Fig. 8), remapping after goal changes, and, by construction, theta sequences (Foster and Wilson, 2007). These representational effects were linked to performance and align with evidence on sparse DG activity and CA3 sequences (Leutgeb and Leutgeb, 2007; Fenton, 2024).
Beyond explaining brain phenomena, our brain-inspired minimal model also proved useful as a module in competitive deep RL agents. The CA3 shift register can be viewed as a sparsely active reservoir that generates finite-length temporal bases, contrasting with the rotational modes of Legendre SSMs and the fading modes of Laguerre SSMs. Its structure resonates with recent shift–diagonal architectures (Fu et al., 2022), but tuned to sparse sensory regimes and navigation tasks. Thus, the model not only explains how theta sequences and place cells form but also provides a normative account of the computational effectiveness of this mechanism for navigation.
The comparison between our hippocampus-inspired model and LSTM agent with dense input demonstrate the distinction of learned strategies and representations, even though the performance indicators are comparable (Figs. A12A11A8A9A10).
Biological Predictions.
Our model predicts that larger environments or sparser inputs require longer sequences for successful navigation, consistent with adaptability of sequence length (Wikenheiser and Redish, 2015; Farooq and Dragoi, 2019). More broadly, it suggests that hippocampal spatial representations could rely largely on intrinsic sequence-generating circuitry, with experience primarily shaping feedforward and readout connections. It also offers an explanation for how place cells can persist despite lesions of entorhinal cortex (Brun et al., 2008; Steffenach et al., 2002; Schlesiger et al., 2015).
Future Directions.
Several extensions follow naturally. On the biological side, incorporating local plasticity rules into the DG–CA3 pathway or CA3 readouts would align the model more closely with known mechanisms while preserving the benefits of prewired dynamics. Interactions with entorhinal inputs and path integration could further clarify their complementary roles in spatial cognition. Moreover, theta sequences have also been observed in other brain areas, raising the possibility of hierarchical coordination across regions.
These ideas resonate with developments in machine learning: structured dynamics akin to SSMs and linear attention are increasingly used in modern sequence models, including LLMs (Fu et al., 2022; Katharopoulos et al., 2020), suggesting that hippocampal-like motifs may illuminate principles of efficient long-range computation. The simplicity of our model also makes it interpretable: sequentially connected CA3 units can be seen as structurally representing trajectories, which could in turn motivate algorithms that minimize internally measured trajectory length without relying on external reward.
On the ML side, combining sparse sequential reservoirs with learned SSMs may yield hybrid architectures capable of adapting flexibly across input regimes. More broadly, a bottom-up approach like ours—focusing on circuit motifs rather than top-down expert informed objectives—suggests a path toward scaling to larger networks and more complex tasks.
RL for neuroscience.
Beyond ML implications, our framework illustrates how modern RL can serve as a testbed for neuroscience theories. By embedding a biologically inspired model within a reinforcement-learning agent, we can study brain computation beyond intermediate objectives such as sensory processing, cognition, or motor control, and consider the entire perception–action cycle—particularly useful for association cortices such as hippocampus and prefrontal cortex where representations reflect both the input and action planning.
Furthermore, the behavioral and representational comparison between CA3 agent and LSTM agent showed a useful direction of research - qualitatively distinct strategies beyond performance benchmarking can provide useful insights for architectural designs.
Conclusion
A minimal, sparsely driven sequence generator in an actor-critic agent not only is sufficient for successful navigation but also for hippocampus-like spatial representations.
Appendix
5 Limitations
Physiological Limitations
Our model has fixed sequential connectivity in CA3, ignoring dynamic processes such as synaptic turnover, plasticity, and sequence length adaptations observed biologically. This simplification may limit the biological realism and generalizability of the model to adaptive neural processes. Additionally, the absence of direct entorhinal cortex inputs means we have not modeled potential interactions between hippocampal sequences and detailed sensory-motor driven activity.
Conceptual Assumptions
We assume sequential activity in CA3 arises exclusively from connectivity rather than sequential inputs, potentially oversimplifying hippocampal dynamics. This assumption excludes possible interactions between external sequential signals and intrinsic CA3 network dynamics, which could affect real-world predictive accuracy.
Reinforcement Learning Methodological Limitations
Our gradient-based reinforcement learning approach lacks clear biological analogs, as organisms likely utilize local Hebbian synaptic updates rather than global gradient propagation. Furthermore, the biological interpretation of the multilayer perceptron decoder remains unclear, posing challenges to direct biological interpretations of our results. Additionally, our use of asynchronous, batch-based training may not accurately reflect the temporal continuity or real-time single-organism learning dynamics seen biologically, thus potentially limiting ecological validity.
Technical Limitations
Our results depend on the environment specifications and hyperparameter settings. We chose a reasonably difficult environment setting where the difference between architectures can be highlighted. It does not imply the effect can generalize to all environments. We tested the agent with L=64 and R=8 on 5 other randomly generated maps, it could converge at a similar speed with the map shown in other results A2. We also tested different frame-skip number, larger frame-skip leads to faster learning, and the effect of sequence length on performance is weaker (Fig. A13). This is probably due to the fact that our environment has sparse obstacles and that larger frame-skip leads to effectively longer sequence and memory. Most results in our report use 4 policies for population based training and 8 environments per worker. These parameters also change the dynamics of learning. having more policies made agents with L=64 learn slower and L=8 learn slower. Further systematic investigation of these factors is required to establish broader applicability.
Taken together, these limitations highlight important areas for further experimental and theoretical investigation.
6 Implementation Details
6.1 Environment
Description Look (dx) Pitch (dy) Strafe Forward Fire Jump Crouch
Forward 0 0 0 1 0 0 0
Strafe Left 0 0 -1 0 0 0 0
Strafe Right 0 0 1 0 0 0 0
Look Left + Forward -20 0 0 1 0 0 0
Look Right + Forward 20 0 0 1 0 0 0
Table A2: Reduced action set used in DeepMind Lab experiments. Each action is defined over the 7-dimensional control space (look, pitch, strafe, forward, fire, jump, crouch).
Since we are mainly interested in modeling the way-finding aspect of navigation [Tolman, 1948, Kuipers, 2000], the MLP also receives average depth of the pixels, down-sampled to 10 horizontal pixels to aid motor control (going around walls and avoiding collisions). This procedure is supposed to mimic the sensory motor collision avoidance habits that do not depend on hippocampus. This depth information is directly routed to the decoder layer, so the representations in the recurrent core and its input are not directly influenced.
6.2 Architecture
Table A3: Network architecture of the CA3 agent. The ResNet encoder is pretrained and frozen during reinforcement learning to isolate the contribution of hippocampal modules.
Component Specification
Visual encoder ResNet using the default parameters from sample factory [Petrenko et al., 2020]
3 convolutional stages (3→16→32→32 channels) each followed by max pooling;
6 residual blocks total (two per stage), each with [ReLU → Conv(3×3) → ReLU → Conv(3×3)] + skip connection;
final ReLU applied to output feature map.
Flatten + MLP head Flatten → Linear(3456→256) → ReLU; output = 256-D embedding.
DG projection Linear(256→16) → BatchNorm1d(momentum=0.05, affine=False) → ReLU; intercept threshold = 2.43.
CA3 (recurrent core) Fixed linear shift-register reservoir; input features = 16, sequence length $L=64$, repetition $R=8$;
total state size $16\times(64+8-1)=1136$.
Decoder MLP: Linear(1136+10→128) → ReLU → Linear(128→128) → ReLU.
Critic head Linear(128→1).
Actor head Linear(128→5) (softmax over 5 discrete actions).
Training scheme Asynchronous PPO [Espeholt et al., 2018, Petrenko et al., 2020] ;
ResNet weights frozen; DG, Decoder, Actor, and Critic trained.
CA3 agent has learnable parameters in DG projection and from CA3 to Decoder, the CA3 output in the full version has 1136*128=149504 parameters. LSTM has learnable parameters $4*(m^{2}+mn)$ where m is hidden size and n is input and output size. Solving this gives a hidden size of roughly 137, which we used in the implementation. SSM agents have matching state size with CA3 agent. They also have block-diagonal recurrent weights, where the weights in each block are obtained through the original implementation by Gu et al. [2020] with zero-order-hold discretizations.
6.3 Actor Critic
We optimize an actor–critic objective with advantage estimates based on the clipped PPO formulation [Schulman et al., 2017]. The total loss combines a clipped policy surrogate, a value regression term, and an entropy regularizer:
$\mathcal{L}(\theta,\phi)=-\mathbb{E}_{t}\!\left[\min\!\bigl(r_{t}(\theta)A_{t},\,\text{clip}(r_{t}(\theta),1-\epsilon,1+\epsilon)A_{t}\bigr)\right]+c_{v}\,\mathbb{E}_{t}[(R_{t}-V_{\phi}(s_{t}))^{2}]-c_{e}\,\mathbb{E}_{t}[\mathcal{H}(\pi_{\theta}(\cdot\mid s_{t}))],$
where $r_{t}(\theta)=\pi_{\theta}(a_{t}\mid s_{t})/\pi_{\theta_{\text{old}}}(a_{t}\mid s_{t})$.
Advantage estimates $A_{t}$ are computed from multi-step discounted returns using generalized advantage estimation (GAE; [Schulman et al., 2015]) with parameters $\gamma$=0.99 and $\lambda$=0.95. Training is performed in the asynchronous actor–learner framework of Sample Factory [Petrenko et al., 2020].
6.4 Training configuration
We trained all agents using the Sample Factory framework [Petrenko et al., 2020] with an asynchronous distributed setup. Each experiment used 32 workers running 8 environments per worker, with experience decorrelated for up to 120 seconds. The environment applied a frame-skip of 8, i.e., each selected action was repeated for 8 frames before the next observation was retrieved.
We used the Asynchronous Proximal Policy Optimization (APPO) algorithm [Petrenko et al., 2020] with discount factor $\gamma=0.99$ and GAE parameter $\lambda=0.95$. Training was performed with a rollout length of 64, a recurrence length of 64, a batch size of 2048, and two batches per epoch. The optimizer used a learning rate of $2\times 10^{-4}$.
The agent architecture followed the design summarized in Table A3. The visual encoder was a ResNet backbone pretrained with an LSTM core on DMLab. For the control experiment, we additionally tested the second convolutional block of ResNet-18 pretrained on ImageNet [He et al., 2015]. The dentate gyrus (DG) module applied batch normalization followed by ReLU activation with an intercept of 2.43, producing 16 sparse feature channels. The decoder consisted of two fully connected layers of 128 units each with ReLU activations.
We employed Population-Based Training (PBT) with four concurrent policies. Policy replacement occurred when the performance gap exceeded 0.05 in relative terms or 0.2 in absolute reward units. Mutations began after 10 million environment steps and were evaluated every 2 million steps. A policy lag tolerance of 35 was used to maintain stability.
Training ran for 108,000 seconds, with checkpoints saved every 120 seconds and milestone checkpoints every 5400 seconds. All experiments were run on CPU, with affinity pinning disabled and input normalization turned off. We used random seeds [1111, 2222, 3333, 4444, 5555]. All other hyperparameters followed the default settings provided in Sample Factory for the DeepMind Lab environment.
6.5 Skaggs’ Spatial Information Measure
We quantify how much information a unit’s activity conveys about the agent’s location in our discrete‐time simulation by a “bits per step” version of Skaggs’ spatial information Skaggs et al. [1992]. Let the environment be divided into $N$ spatial bins, and define:
  • $p_{i}$ — fraction of timesteps (steps) spent in bin $i$ (occupancy probability),
  • $\lambda_{i}$ — mean activity rate in bin $i$, measured in activity per step,
  • $\lambda=\sum_{i=1}^{N}p_{i}\,\lambda_{i}$ — overall activity rate (activity per step).
First, the information conveyed per timestep is
$I_{\mathrm{step}}\;=\;\sum_{i=1}^{N}p_{i}\,\lambda_{i}\log_{2}\!\biggl(\frac{\lambda_{i}}{\lambda}\biggr)\quad\text{[bits/step].}$
This measures the average reduction in positional uncertainty (in bits) each simulation step provides.
Interpretation:
  • Uniform firing ($\lambda_{i}=\lambda$ for all $i$) yields zero information ($I_{\mathrm{step}}=I_{\mathrm{spike}}=0$).
  • Elevated $\lambda_{i}$ in particular bins gives positive contributions proportional to $p_{i}\,\lambda_{i}$ (for bits/step) or $p_{i}\,(\lambda_{i}/\lambda)$ (for bits/spike).
  • Bins visited rarely (small $p_{i}$) contribute less, guarding against over‐weighing seldom‐visited locations.
In our simulations, $\lambda_{i}$ and $\lambda$ are estimated from spatially binned sum of activity divided by the number of steps in each bin. Bins with $\lambda_{i}=0$ are omitted (treating $0\cdot\log 0=0$).
7 Supplementary Figures
Figure A1: Robustness to changes of visual encoder. Training performance of CA3 agent with the ResNet layer 2 output as visual encoder. Random seeds: [1111, 2222, 3333, 4444, 5555].
Figure A2: Robustness to changes of realization of environment. Training performance on 5 other randomly generated maps. Lines with the same color are results from random seeds [0, 42]
Figure A3: Comparison to state-space model (SSM) based architectures. Training performance of agents with fixed SSM cores described by Gu et al. [2020]. legt: Legendre bases with fixed memory horizon. legs: Legendre bases with infinite memory from the beginning of an episode. lagt: Laguerre bases, i.e. memory with decay. All three SSMs were implemented with the same state size as our CA3 module, each input feature is expanded into 64+8-1=71 states. The SSMs were fixed, with three additional learnable parameter controlling the input, recurrent scales and timestep length for discretizations, as a common practice described by Gu et al. [2021].
Figure A4: More examples of rate maps from decoder layer 1 in the last training epoch. Showing the random 16 units out of the 50 units with the best SI. Unit id displayed on top.
Figure A5: More examples of rate maps from decoder layer 2 in the last training epoch. Showing the random 16 units out of the 50 units with the best SI. Unit id displayed on top.
Figure A6: Distributions of spatial information (SI) rates becomes heavy tailed over learning. SI rate distributions for different layers (brain regions rows) over the progress of learning (colors) and change of reward site (purple). Y axes: unit fraction density.
Figure A7: Population representations stabilize over learning and remap. PV correlations across training. Although performance saturates by epochs 5–6, representational changes continue. The last epoch shows the effect of learning a new goal location.
Figure A8: Vision centered behavioral strategy in LSTM-based agents. Occupancy map and mean movement direction (white) of LSTM-based agent’s trajectories across 5 epochs of training.
Figure A9: Example ratemap of LSTM output units. Place fields from the LSTM core (randomly selected units). Each panel shows the normalized spatial activity of one unit.
Figure A10: LSTM-based agent exhibits less input orthogonalization. Population vector correlation of LSTM inputs across epochs. Color indicates Pearson correlation; rows show epochs 0–4.
Figure A11: LSTM-based agent exhibits less input sparseness. Learned dense input features in LSTM agent. Each tile shows the spatial receptive field of one feature.
Figure A12: No clear distance-dependent representation of LSTM-based agents. A Displacement-dependent Pearson correlation kernel of population vectors. Axes show spatial shifts $(\Delta x,\Delta y)$; values are averaged correlations. Rows correspond to different layers (brain regions) of the network, columns to learning stages (epochs). B Left: summary of population vector correlation as a function of distance. Right: summary of population vector correlation as a function of orientation. Radial coordinate range: (-0.5,1.1).
Figure A13: Robustness to changes in training parameters. Training performance with different environment action repeats and number of policies in population based training. Left: environment frame skip / action repeat = 8. Right: environment frame skip / action repeat = 4. The frame skip controls the fine-graininess of actions, which is not required for the current navigation task. On the other hand, larger frame skip effectively make sequences propagate over longer traversals.